# AutoLibra: Agent Metric Induction from Open-Ended Feedback

### v1

@@ Abstract
Agents are predominantly evaluated and optimized via task success metrics, which are coarse, rely on manual design from experts, and fail to reward intermediate emergent behaviors. We propose AutoLibra [Uncaptioned image], a framework for agent evaluation, that transforms open-ended human feedback e.g.If you find that the button is disabled, don’t click it again”, or “This agent has too much autonomy to decide what to do on its own” into metrics for evaluating fine-grained behaviors in agent trajectories. AutoLibra accomplishes this by grounding feedback to an agent’s behavior, clustering similar positive and negative behaviors, and creating concrete metrics with clear definitions and concrete examples, which can be used for prompting LLM-as-a-Judge as evaluators. We further propose two meta-metrics to evaluate the alignment of a set of (induced) metrics with open feedback: “coverage” and “redundancy”. Through optimizing these meta-metrics, we experimentally demonstrate AutoLibra’s ability to induce more concrete agent evaluation metrics than the ones proposed in previous agent evaluation benchmarks and discover new metrics to analyze agents. We also present two applications of AutoLibra in agent improvement: First, we show that AutoLibra-induced metrics serve as better prompt-engineering targets than the task success rate on a wide range of text game tasks, improving agent performance over baseline by a mean of 20%. Second, we show that AutoLibra can iteratively select high-quality fine-tuning data for web navigation agents. Our results suggest that AutoLibra is a powerful task-agnostic tool for evaluating and improving language agents.
@@ 1 Introduction
Humans readily acquire skills from open-ended instructions and feedback from others (Tomasello et al., 1993). These instructions and feedback are internalized for self-regulated learning (Pintrich & Zusho, 2002; Nicol & Macfarlane-Dick, 2006), providing internal signals for continuous improvement. Drawing inspiration from this process, we investigate how well AI agents can benefit from open-ended human feedback through induction of generalizable metrics.
We can characterize existing evaluations and rewards for AI agents into two main classes: (1) goal-oriented evaluation – whether the agents have fulfilled the given task, e.g. benchmarks (Zhou et al., 2024a; Jimenez et al., 2024; Chan et al., 2024; Paglieri et al., 2024) and reward modeling approaches (Pan et al., 2024; Chen et al., 2025b; Choudhury, 2025) and (2) behavior evaluation – the manner the agents perform the task, e.g. social agent and human-agent interaction benchmarks (Zhou et al., 2024b; Shao et al., 2024) and agent failure mode analysis (Pan et al., 2025; Zhang et al., 2023; Yang et al., 2023). [^1] Note that the difference between the two classes is not whether the evaluation is based on the outcome or the process. Both evaluations can be performed at every step or in the final step. Goal-oriented evaluation is often designed to be verifiable, but is not sufficiently fine-grained or comprehensive to diagnose agents’ behavior problems or locate improvement bottlenecks (Yehudai et al., 2025). Although the above can be complemented with behavior evaluation, it requires manual design of metrics based on top-down heuristics (Zhou et al., 2024b), or thematic analysis of the agent’s behavior (Shao et al., 2024; Pan et al., 2025). This manual design process requires the labor of experts in annotations and classification, which limits its scalability.
In this paper, we introduce AutoLibra [Uncaptioned image], a metric induction method, as a novel agent evaluation framework that mitigates the limitations of current evaluation paradigms. AutoLibra is an evaluation tool that induces interpretable metrics for AI agents from open-ended human feedback, which can be collected from end users of AI agents or experts. This offers two key advantages: (1) It is much easier to provide concrete feedback for trajectories than creating metrics, and (2) AutoLibra allows us to evaluate agents from the perspective of the users. AutoLibra-induced metrics provide concrete definitions of behaviors that the model-based evaluation method should look for when evaluating AI agents, which could be used to understand agent behavior, as well as optimization targets to improve agents.
The AutoLibra evaluation process is designed to provide a closed-loop feedback signal for the induction process. The agent trajectories used in the induction process are scored by LLM-as-a-Judge (Zheng et al., 2023) on the induced metrics. The evaluation process (§2.3) then tries to match the feedback aspects, e.g. “recipe does not contain quinoa”, with the traits, e.g. task-requirement-achievement. In this way, we can meta-evaluate the quality of the metrics: (i) coverage (what proportion of feedback aspects can be matched with an agent trait), and (ii) redundancy of the metrics (what proportion of the detected traits are not mentioned by humans). These two metrics provide an overall statistical picture of the quality of the induced metrics. Based on these two metrics, we can search for the set of metrics with the lowest redundancy within those with the highest coverage. As shown in §3.1, we find that as the number of metrics increases, the redundancy increases, and the coverage ultimately converges to the maximum coverage.
With AutoLibra, our aim is to answer the following research questions:
RQ1:
How aligned are the results from AutoLibra for each step with human judgment?
RQ2:
Can AutoLibra provide more insight into agent behavior that were not provided by expert designed metrics?
RQ3:
Can AutoLibra provide optimization signals for human agent engineers and agent learning algorithms to improve agents’ performance?
Key findings: Experiments within multiple agent domains, including collaborative agents (Shao et al., 2024), social agents (Zhou et al., 2024b), web agents (Zhou et al., 2024a; He et al., 2024), and text game agents (Paglieri et al., 2024; Cloos et al., 2024), demonstrate that AutoLibra is able to induce fine-grained and interpretable metrics with high coverage and low redundancy in unseen human feedback with 80 trajectories annotated with one feedback for each trajectory per dataset. These metrics are more concrete, and some of them were even overlooked in expert designed metrics or error analysis (§4). The fact that AutoLibra can iteratively discover new, emergent metrics (§3.2) throughout the agent optimization process, and provide salient and human-interpretable optimization signals helps improve the performance in Baba-Is-AI by over 20% (§5.1), and web-browsing task WebVoyager by 5% on top of previous model-based rejection fine-tuning methods (§5.2) with only 18 trajectories annotated with one feedback for each trajectory in each improvement iteration.
@@ 2 AutoLibra [Uncaptioned image]
To address the limitations of existing evaluation paradigms, AutoLibra [Uncaptioned image] is designed to meet the following desiderata: (1) induced from agent behavior: This ensures that metrics are grounded in agent trajectories rather than predefined by human experts, (2) self-validating: Allows choosing minimal set of metrics that cover unseen human feedback with sufficient abstraction to be useful across different tasks, and (3) generalizable: Applicable to various agent environments, independent of domain-specific design.
Based on feedback data collected from humans (§2.1), AutoLibra achieves these desiderata through a closed-loop pipeline consisting of two processes: Induction Process that converts agent behaviors and corresponding feedback into metrics, (§2.2) and Evaluation Process that predicts ratings and quality of new agent behaviors on the induced metrics (§2.3).
@@ 2.1 Collecting human feedback
In this paper, we use human feedback from two groups: (1) End-users – for agents that interact directly with humans, we use the feedback from the users who interact and converse with the agents. CoGym (Shao et al., 2024) is the environment that belongs to this category, and we use the user comments collected in their study, resulting in 197 trajectories with feedback. (2) Experts – for agents that do not directly interact with humans, we use the feedback from human annotators (five authors in this paper) who observe agent trajectories. All other environments belong to this category, these being Sotopia (Zhou et al., 2024b), WebArena (Zhou et al., 2024a), WebVoyager (He et al., 2024), Baba-is-ai (Cloos et al., 2024), and MiniHack (Samvelyan et al., 2021). For each trajectory, we collect only one element of feedback; feedback is given based on the complete agent trajectories. [^2] While in theory we can leverage feedback on specific steps to achieve better feedback grounding and multiple feedback for single trajectory, we leave it as future work.
Figure 2: Feedback Grounding
@@
Annotators are instructed to explicitly indicate the aspects of agent behavior that they classify as good or bad, and to avoid general comments such as "The agent is good at solving the task". The annotators can also choose from a TTY (TeleTYpewriter) or a web interface; in both cases the annotator is provided with the agent’s task and then view the agent’s observation and actions step by step, in text form.  [^3] While viewing screenshots is standard for web navigation tasks, we keep the observation format consistent across agents and humans to encourage more grounded feedback. For multi-agent tasks, we annotate each agent’s trajectory in a given interaction separately. For Sotopia (Zhou et al., 2024b), WebArena (Zhou et al., 2024a), and WebVoyager (He et al., 2024), we annotate 100 trajectories of agents based on GPT-4 (Achiam et al., 2023) with feedback for each dataset. For experiments in §5 we annotate 18 trajectories for each dataset in each iteration. The annotation process is fast: Human annotators spend less than 5 minutes to provide feedback for each trajectory; §4, we randomly hold out 20% of the trajectories for validation.
2.2 Induction Process: agent trajectories and human feedback $\rightarrow$ evaluation metrics
Feedback Grounding
[…]
@@ 2.2 Induction Process: agent trajectories and human feedback $\rightarrow$ evaluation metrics
Illustrated in Fig. 2, in this step, we feed the trajectory and the feedback into the LLM (we use GPT-4o (OpenAI et al., 2024) as it yields good results in our pilot experiments) and prompt the LLM with the following instructions: (1) break down the feedback into bullet points; (2) for each bullet point, find the corresponding part of the trajectory to which the feedback refers. Finally, we use constrained decoding to force GPT-4o to output the aspects in the previous format. In our experiments, we find that on most datasets, for each trajectory, the LLM can generate one to five aspects, with a mean of one to two aspects.
@@
However, clustering similar agent behaviors together is challenging for statistical clustering methods.444 In preliminary experiments, we tried to use K-means clustering on the aspect vectors generated by text-embedding-3-large, but the clusters are mostly based on tasks and not on the behaviors. Inspired by LLM-based semantic clustering and concept induction methods Viswanathan et al. (2024); Lam et al. (2024), we prompt an LLM (o3-mini high [^5] https://openai.com/index/openai-o3-mini/, as it produces the most accurate coverage and redundancy scores as evaluated later) to cluster the aspects into metrics. As illustrated in Fig. 3, we gather all the aspects of $M$ trajectories and cluster into $N$ metrics, where $N$ is a parameter set through the optimization process (§3.1). We provide the LLM with the following instructions: The granularity of the grouping should be minimal; only very similar behaviors are grouped together; but don’t limit to one particular website or one particular character, which empirically makes the metrics more concrete but still applicable across different tasks.
2.3 Evaluation Process: evaluating agents and the quality of the induced metrics
@@
Figure 4: Evaluating agents with
induced metrics
@@
The last component of the loop is the meta-evaluation, i.e. evaluating the evaluation metrics induced by AutoLibra. This step matches the traits detected by the LLM-as-a-Judge with aspects grounded from the human feedback. The goal is to verify whether (1) the induced metrics cover the behaviors the human annotators care about, and (2) LLM-as-a-Judge can produce accurate evaluation results based on the induced metrics. In the previous example, if the respond-promptly is extracted as a metric, and the LLM-as-a-Judge has the same opinion as the human annotators, then this aspect is considered as successfully covered. If either a similar metric was not extracted, or the LLM-as-a-Judge assigns a different score, then this aspect is considered as not covered.
Figure 5: Meta evaluation
@@ 3.1 Metric Optimization
As illustrated in Fig. 6, we optimize the metric induction process to maximize coverage and minimize redundancy. Among the two, we prioritize coverage of the metrics to provide a comprehensive evaluation of the agent behavior, while minimizing overlap within the metrics to avoid redundancy, thus maximizing the utility of induced metrics. To optimize for this objective, we generate 20 different metrics, with $N$ ranging from 4 to 13, and calculate the coverage and redundancy of the metrics in human feedback.
@@
We then select metrics with a coverage of at least the highest coverage minus 1%, and the lowest redundancy. This is performed iteratively, by resetting the range of $N$ to the number of metrics selected previously $\pm$2, repeating until the coverage and redundancy of the selected metrics converge, normally within 3 iterations. While this optimization process is simple, experiments with various other optimization strategies, including genetic algorithms and iterative clustering saw none of them yield better results than the simple strategy. Fig. 7 shows the highest coverages of the metrics of size $N$, which converge around $N=6$ to $10$ depending on the datasets. The best coverage on Sotopia (Zhou et al., 2024b) is the lowest among all four datasets, $60\%$, likely due to the diversity of the tasks in the dataset, while coverage on WebArena (Zhou et al., 2024a) and WebVoyager (He et al., 2024) are the highest, $88\%$. We also find that the coverage of the held-out trajectories is only slightly worse ($<5\%$) than the trajectories we use to induce the metrics, which is expected since we use the exact examples extracted from the latter. Lastly, we show that the good and bad behaviors are crucial in the metrics, dropping which resulting in up to $30\%$ coverage decrease (on CoGym (Shao et al., 2024)).
@@ 3.2 Iterative Metric Induction
When applying AutoLibra to agent optimization, we can iteratively induce new metrics, as agents develop new failure modes or new behaviors as they improve, which is useful for tracking agents’ progress across different iterations. [^6] Alternatively, a new set of metrics can be induced from scratch for each iteration - in practice, we do not find that this results in any coverage loss, but we choose the former method for consistency
To do this, we modify the behavior clustering step, by providing the LLM with the existing metrics and their definitions, and ask the LLM not to change the definitions of the existing metrics, to only add new behaviors to the existing metrics, and add new metrics if necessary. We apply the same optimization strategy as in the metric optimization step ensure the newly induced metrics cover emerging behaviors and do not overlap with existing metrics.

### v3

# AutoLibra: Agent Metric Induction from Open-Ended Human Feedback

@@ Abstract
Agents are predominantly evaluated and optimized via task success metrics, which are coarse, rely on manual design from experts, and fail to reward intermediate emergent behaviors. We propose AutoLibra [Uncaptioned image] , a framework for agent evaluation, that transforms open-ended human feedback e.g.If you find that the button is disabled, don’t click it again”, or “This agent has too much autonomy to decide what to do on its own” into metrics for evaluating fine-grained behaviors in agent trajectories. AutoLibra accomplishes this by grounding feedback to an agent’s behavior, clustering similar positive and negative behaviors, and creating concrete metrics with clear definitions and concrete examples, which can be used for prompting LLM-as-a-Judge as evaluators. We further propose two meta-metrics to evaluate the alignment of a set of (induced) metrics with open feedback: “coverage” and “redundancy”. Through optimizing these meta-metrics, we experimentally demonstrate AutoLibra’s ability to induce more concrete agent evaluation metrics than the ones proposed in previous agent evaluation benchmarks and discover new metrics to analyze agents. We also present two applications of AutoLibra in agent improvement: First, we show that AutoLibra serve human prompt engineers for diagonalize agent failures and improve prompts iterative. Moreover, we find that AutoLibra can induce metrics for automatic optimization for agents, which makes agents improve through self-regulation. Our results suggest that AutoLibra is a powerful task-agnostic tool for evaluating and improving language agents.
@@ 1 Introduction
Humans readily acquire skills from open-ended instructions and feedback from others (Tomasello et al., 1993). These instructions and feedback are internalized for self-regulated learning (Pintrich & Zusho, 2002; Nicol & Macfarlane‐Dick, 2006), providing internal signals for continuous improvement. Drawing inspiration from this process, we investigate how well AI agents can benefit from open-ended human feedback through induction of generalizable metrics.
In this paper, we introduce AutoLibra [Uncaptioned image] , a metric induction method, as a novel agent evaluation framework that mitigates the limitations of current evaluation paradigms. AutoLibra is an evaluation tool that induces interpretable metrics for AI agents from open-ended human feedback, which can be collected from end users of AI agents or experts. This offers two advantages: (1) It is much easier to provide concrete feedback for trajectories than creating metrics, and (2) AutoLibra allows us to evaluate agents from the perspective of the users. AutoLibra-induced metrics provide concrete definitions of behaviors that the model-based evaluation method should look for, which could be used to understand agent behavior, as well as optimization targets to improve agents.
The AutoLibra evaluation process is designed to provide a closed-loop feedback signal for the induction process. The agent trajectories used in the induction process are scored by LLM-as-a-Judge (Zheng et al., 2023) on the induced metrics. The evaluation process (§2.3) then tries to match the feedback aspects, e.g. “recipe does not contain quinoa”, with the traits, e.g. task-requirement-
achievement
. In this way, we can meta-evaluate the quality of the metrics: (i) coverage (what proportion of feedback aspects can be matched with an agent trait), and (ii) redundancy of the metrics (what proportion of the detected traits are not mentioned by humans). These two metrics provide an overall statistical picture of the quality of the induced metrics. Based on these two metrics, we can search for the set of metrics with the lowest redundancy and the highest coverage. As shown in §3.1, we find that as the number of metrics increases, the redundancy increases, and the coverage ultimately converges to the maximum coverage. With AutoLibra, our aim is to answer the following research questions:
  • RQ1:
    How well do AutoLibra’s step-wise results align with human judgment?
  • RQ2:
    Does AutoLibra provide insights into agent behavior beyond expert-designed metrics?
  • RQ3:
    Can AutoLibra provide optimization signals for improving agents’ performance?
Experiments within multiple agent domains, including collaborative agents (Shao et al., 2024), social agents (Zhou et al., 2024b), web agents (Zhou et al., 2024a; He et al., 2024), and text game agents (Paglieri et al., 2024; Cloos et al., 2024), demonstrate that AutoLibra is able to induce fine-grained and interpretable metrics with high coverage and low redundancy in unseen human feedback with 80 trajectories per dataset annotated with one feedback for each. These metrics are more concrete, and some of them were even overlooked in expert designed metrics or error analysis (§4). AutoLibra can iteratively discover new, emergent metrics (§3.2) throughout the agent optimization process, and provide optimization signals helps improve the performance of frontier LLM in a challenging 2D text game by over 20% (§5) in 3 stages with only 18 trajectory annotated per stage.
@@ 2 AutoLibra [Uncaptioned image]
To address the limitations of existing evaluation paradigms, AutoLibra [Uncaptioned image] is designed to meet the following desiderata: (1) induced from agent behavior: This ensures that metrics are grounded in agent trajectories rather than predefined by human experts, (2) self-validating: Allows choosing minimal set of metrics that cover unseen human feedback with sufficient abstraction to be useful across different tasks, and (3) generalizable: Applicable to various agent environments, independent of domain-specific design. Based on feedback data collected from humans (§2.1), AutoLibra achieves these desiderata through a closed-loop pipeline consisting of two processes: Induction Process that converts agent behaviors and corresponding feedback into metrics, (§2.2) and Evaluation Process that predicts ratings and quality of new agent behaviors on the induced metrics (§2.3).
@@ 2.1 Collecting human feedback
In this paper, we use human feedback from two groups: (1) End-users – for agents that interact directly with humans, we use the feedback from the users who interact and converse with the agents. CoGym (Shao et al., 2024) is the environment that belongs to this category, and we use the user comments collected in their study, resulting in 197 trajectories with feedback. (2) Experts – for agents that do not directly interact with humans, we use the feedback from human annotators (five authors in this paper) who observe agent trajectories. All other environments belong to this category, these being Sotopia (Zhou et al., 2024b), WebArena (Zhou et al., 2024a), WebVoyager (He et al., 2024), Baba-is-ai (Cloos et al., 2024), and MiniHack (Samvelyan et al., 2021). For each trajectory, we collect only one element of feedback based on the complete agent trajectories. [^1] While in theory we can leverage feedback on specific steps to achieve better feedback grounding and multiple feedback for single trajectory, we leave it as future work.
@@
Annotators are instructed to explicitly indicate the aspects of agent behavior that they classify as good or bad, and to avoid general comments such as "The agent is good at solving the task". The annotators can also choose from a terminal or a web interface; in both cases the annotator is provided with the agent’s task and then view the agent’s observation and actions step by step, in text form. [^2] While viewing screenshots is standard for web navigation tasks, we keep the observation format consistent across agents and humans to encourage more grounded feedback. For multi-agent tasks, we annotate each agent’s trajectory in a given interaction separately. For Sotopia (Zhou et al., 2024b), WebArena (Zhou et al., 2024a), and WebVoyager (He et al., 2024), we annotate 100 trajectories of agents based on GPT-4 (Achiam et al., 2023) with feedback for each dataset. For experiments in §5 we annotate 18 trajectories for each dataset in each iteration. The annotation process is fast: Human annotators spend less than 5 minutes to provide feedback for each trajectory; §4, we randomly hold out 20% of the trajectories for validation.
2.2 Induction Process
Feedback Grounding The feedback of human annotators can contain multiple aspects; e.g. “AI agent was pretty good at giving me a consistent itinerary and vacation plan, although it froze on the last couple of minutes.”, collected from human annotators in CoGym (Shao et al., 2024), contains a positive aspect about the agent’s ability to generate a consistent itinerary, and a negative aspect about the agent freezing at the end. Here we define an aspect as a triple $(\texttt{behavior},\texttt{feedback},\texttt{sign})$. In the positive aspect of the previous example, the behavior is the agent’s actions to create a 20-day itinerary for the Maldives, the feedback is that the created itinerary is consistent and the sign is positive. This grounding procedure is similar to the coding procedure in thematic analysis.
@@ 2.2 Induction Process
We feed the trajectory and the feedback into the LLM (we use GPT-4o (OpenAI et al., 2024) as it yields good results in our pilot experiments) and prompt the LLM with the following instructions: (1) break down the feedback into bullet points; (2) for each bullet point, find the corresponding part of the trajectory to which the feedback refers. Finally, we use constrained decoding to force GPT-4o to output the aspects in the previous format. In our experiments, we find that on most datasets, for each trajectory, the LLM can generate one to five aspects, with a mean of one to two aspects.
@@
However, clustering similar agent behaviors together is challenging for statistical clustering methods. [^3] In preliminary experiments, we tried to use K-means clustering on the aspect vectors generated by embedding model text-embedding-3-large, but the clusters are mostly based on tasks and not on the behaviors. Inspired by LLM-based semantic clustering and concept induction methods (Viswanathan et al., 2024; Lam et al., 2024), we prompt an LLM (o3-mini high [^4] https://openai.com/index/openai-o3-mini/, as it produces the most accurate coverage and redundancy scores as evaluated later) to cluster the aspects into metrics. As illustrated in Fig. 7, we gather all the aspects of $M$ trajectories and cluster into $N$ metrics, where $N$ is a parameter set through the optimization process (§3.1). We provide the LLM with the following instructions: The granularity of the grouping should be minimal; only very similar behaviors are grouped together; but don’t limit to one particular website or one particular character, which empirically makes the metrics more concrete but still applicable across different tasks.
2.3 Evaluation Process
@@
The final loop component is the meta-evaluation, i.e. evaluating the evaluation metrics induced by AutoLibra. This step matches the traits detected by the LLM-as-a-Judge with aspects grounded from the human feedback. The goal is to verify whether (1) the induced metrics cover the behaviors the human annotators care about, and (2) LLM-as-a-Judge can produce accurate evaluation results based on the induced metrics. In the previous example, if the respond-promptly is extracted as a metric, and the LLM-as-a-Judge has the same opinion as the human annotators, then this aspect is considered as successfully covered. If either a similar metric was not extracted, or the LLM-as-a-Judge assigns a different score, then this aspect is considered as not covered.
@@ 3.1 Metric Optimization
Illustrated in Fig. 2, we optimize the metric induction process to maximize coverage and minimize redundancy. Among the two, we prioritize coverage of the metrics to provide a comprehensive evaluation of the agent behavior, while minimizing overlap within the metrics to avoid redundancy, thus maximizing the utility of induced metrics. To optimize for this objective, we generate 20 different sets of metrics, with metric count $N$ ranging from 4 to 13, and calculate the coverage and redundancy of the metrics in human feedback. We then select metrics with a coverage of at least the highest coverage minus 1%, and the lowest redundancy. This is performed iteratively, by resetting the range of $N$ to the number of metrics selected previously $\pm$2, repeating until the coverage and redundancy of the selected metrics converge, normally within 3 iterations. While this optimization process is simple, experiments with various other
@@
optimization strategies, including genetic algorithms and iterative clustering saw none of them yield better results than the simple strategy. Fig. 3 shows the highest coverages of the metrics of size $N$, which converge around $N=6$ to $10$ depending on the datasets. The best coverage on Sotopia (Zhou et al., 2024b) is the lowest among all four datasets, $60\%$, likely due to the diversity of the tasks in the dataset, while coverage on WebArena (Zhou et al., 2024a) and WebVoyager (He et al., 2024) are the highest, $88\%$. We also find that the coverage of the held-out trajectories is only slightly worse ($5\%$) than the trajectories we use to induce the metrics, which is expected since we use the exact examples extracted from the latter. Lastly, we show that the good and bad behaviors are crucial in the metrics, dropping which resulting in up to $30\%$ coverage decrease on CoGym.
@@ 3.2 Iterative Metric Induction
When applying AutoLibra to agent optimization, we can iteratively induce new metrics, as agents develop new failure modes or new behaviors as they improve, which is useful for tracking agents’ progress across different iterations. [^5] Alternatively, a new set of metrics can be induced from scratch for each iteration - in practice, we do not find that this results in any coverage loss, but we choose the former method for consistency To do this, we modify the behavior clustering step: we provide the LLM with the existing metrics and definitions, asking the LLM not to change the definitions of the existing metrics, to only add new behaviors to the existing metrics, and to add new metrics if necessary. We apply the same optimization strategy as in the metric optimization step ensure the newly induced metrics cover emerging behaviors and do not overlap with existing metrics.